suppose you just have enough dimes and quarters to pay for a loaf of bread and a quart of milk, which cost $3.45. you have a total of 15 dimes and quarters.
let d = the number of dimes you have
let q = the number of quarters you have
a. write a system of equations that models the information given above.
b. solve the system for the number of dimes and quarters you have.

Answer by checkley75(3666)

D+Q=15 D=15-Q NOW SUBSTITUTE (15-Q) FOR D IN THE SECOND EQUATION
10D+25Q=345
10(15-Q)+25Q=345
150-10Q+25Q=345
15Q=345-150
15Q=195
Q=195/15
Q=13 NUMBER OF QUARTERS.
D+13=15
D=15-13
D=2 NUMBER OF DIMES
PROOF
10*2+25*13=345
20+325=345
345=345